OpenStudy (anonymous):

how to graph y=2x-13?? with medal

OpenStudy (ranga):

This is the equation for a straight line. The quickest way to graph it is to find the x and y intercepts. x-intercept is when y = 0 and y-intercept is when x = 0

OpenStudy (anonymous):

how can you show me how to graph

OpenStudy (ranga):

y = 2x - 13 y-intercept is when x = 0 y = 0 - 13 = -13 x-intercept is when y = 0 0 = 2x - 13 x = 13/2 = 6.5 Mark off 6.5 on the x axis Mark off -13 on the y-axis. Using a ruler join draw a straight line passing through those two points.

OpenStudy (anonymous):

i mean how to graph y=2x-13?? graph only no need to solve

OpenStudy (anonymous):

because the question is graph the line through (5.-3) with m=2

OpenStudy (anonymous):

so the anwer is y=2x-13 how to graph??

OpenStudy (ranga):



OpenStudy (ranga):

To graph the straight line you just need two points. And I have shown you a way to quickly get two points on the straight line.